OpenStudy (anonymous):

f(x)=x-5/x+1 g(x) x+2/x-3 fog=? gof=? fof=? gog=?

OpenStudy (anonymous):

It's basically say f of g (x) which means f(g(x)) meaning you substitute g(x) function in f(x). \[\huge f(g(x)) = \frac{ \frac{ x+2 }{ x-3 }-5 }{ \frac{ x+2 }{ x-3 }+1 }\] Put g(x) every place there is a x in function f(x).

OpenStudy (anonymous):

Things will get cancelled so don't get worried, so use this same method and try it with others.

OpenStudy (anonymous):

Thank you!